Please Graph equality. Thanks

2x + 5y ≥ 10

Answer by haileytucki(390)

2x+5y=10
To find the x-intercept, substitute in 0 for y and solve for x.
2x+5(0)=10
Multiply 5 by each term inside the parentheses.
2x+0=10
Remove the 0 from the polynomial; adding or subtracting 0 does not change the value of the expression.
2x=10
Divide each term in the equation by 2.
(2x)/(2)=(10)/(2)
Simplify the left-hand side of the equation by canceling the common factors.
x=(10)/(2)
Simplify the right-hand side of the equation by simplifying each term.
x=5
To find the y-intercept, substitute in 0 for x and solve for y.
2(0)+5y=10
Multiply 2 by each term inside the parentheses.
0+5y=10
Combine all similar expressions.
5y=10
Divide each term in the equation by 5.
(5y)/(5)=(10)/(5)
Simplify the left-hand side of the equation by canceling the common factors.
y=(10)/(5)
Solve the equation.
y=2
These are the x and y intercepts of the equation 2x+5y=10.
x=5, y=2

points on your graph will be (-5,3), (5,0) , (0,2)...the line will run directly through all given points as if it were infinite